### v1

@@ 7 Conclusion
Modeling high-dimensional tensor time series has gathered increasing attention recently, owing to the availability of multidimensional datasets beyond the classical panel data structure. This paper considers matrix and tensor factor models with a CP low-rank structure, offering a generalization of classical vector factor models. We develop iterative simultaneous orthogonalization estimation procedures based on contemporary covariance, preserving the tensor data structure. Theoretical properties such as the rate of convergence and limiting distributions are investigated, assuming each tensor dimension is comparable to or greater than the number of observations, and the tensor rank might be fixed or divergent.
In contrast to auto-covariance-based estimation methods, we explore information from contemporary data and are also able to consistently estimate loadings and factors for uncorrelated tensor observations where auto-covariance methods might fail. Additionally, we propose two generalized eigenvalue-ratio estimators for rank selection and justify their consistency. A comprehensive simulation study underscores the merits of our proposed method compared to existing methods. Furthermore, empirical applications regarding sorted portfolios and international trade flows showcase the practical relevance of our approach.
@@ Appendix A Proofs of Main Theorem
Part (i). Recall $\widehat{\Sigma}=T^{-1}\sum_{t=1}^{T}{\cal Y}_{t}\otimes{\cal Y}_{t}$, $\Theta=\mathbb{E}F_{t}F_{t}^{\top}$, $a_{i}=\mbox{vec}(a_{i1}\otimes a_{i2}\otimes\cdots\otimes a_{iK})$, $d=d_{1}d_{2}...d_{K}$. Let $e_{t}=\mbox{vec}({\cal E}_{t})$. Write
@@
Part (ii). For simplicity, consider the most extreme case where $\min\{\lambda_{i}-\lambda_{i+1},\lambda_{i}-\lambda_{i-1}\}\leq c\lambda_{r}$ for all $i$, with $\lambda_{0}=\infty,\lambda_{r+1}=0$, and $c$ is sufficiently small constant. In such cases, we need to employ Procedure 2 to the entire sample covariance tensor $\widehat{\Sigma}$. Let the eigenvalue ratio $w:=\lambda_{1}/\lambda_{r}=O(r)$. Without loss of generality, assume $\Theta_{11}\geq\Theta_{22}\geq\cdots\geq\Theta_{rr}$. In general, the statement in the theorem holds for number of initialization $L\geq Cd^{2}\vee Cdr^{2w^{2}}$, where $a\vee b=\max\{a,b\}$. We prove the statements through induction on factor index $i$ starting from $i=1$ proceeding to $i=r$. By the induction hypothesis, we already have estimators such that
@@
and $R^{(0)}=\phi^{(0)}$ is defined in (19) in Theorem 4.1.
@@
$\geq\left\|\Theta_{ii}\prod_{k=1}^{K}(a_{ik}^{\top}\widetilde{a}_{\ell k})^{2}\right\|_{2}-\left\|\sum_{(j_{1},j_{2})\neq(i,i)}^{r}\Theta_{j_{1},j_{2}}\prod_{k=1}^{K}(a_{j_{1}k}^{\top}\widetilde{a}_{\ell k})(a_{j_{2}k}^{\top}\widetilde{a}_{\ell k})\right\|_{2}-\left\|\Psi\times_{k=1}^{2K}\widetilde{a}_{\ell k}\right\|_{2}$
where $\Psi$ is defined by unfolding $\Psi^{*}$ into a $d_{1}\times d_{2}\times\cdots\times d_{K}\times d_{1}\times d_{2}\times\cdots\times d_{K}$ tensor. Let $\psi_{i}=CR^{(0)}/\Theta_{ii}+\delta_{1}^{2}w$. By (A) and the last part of the proof of Lemma A.2, as $\|\Psi^{*}\|_{2}/\lambda_{r}\leq\phi_{0}^{2}$ and $(1+\delta_{1})\prod_{k=2}^{K}(\delta_{k}+\psi_{i})w\leq\phi_{0}^{2}$, it follows that
@@
$\|\Psi^{*}\|_{\rm 2}$ $\leq C\lambda_{1}\left(\sqrt{\frac{r+\log T}{T}}+\frac{(r+\log T)^{1/\gamma}}{T}\right)+C\left(\sqrt{\frac{d\log(d)}{T}}+\frac{d\log(d)}{T}+1\right)$
$\qquad\quad+C\lambda_{1}^{1/2}\left(\sqrt{\frac{d\log(d)}{T}}+\frac{d\log(d)}{T}\right).$ (38)
@@
Let $\Upsilon_{0}=T^{-1}\sum_{t=1}^{T}\sum_{i,j=1}^{r}f_{it}f_{jt}a_{i}a_{j}^{\top}$, $\overline{\mathbb{E}}(\cdot)=\mathbb{E}(\cdot|f_{it},1\leq i\leq r,1\leq t\leq T)$. Define $e_{t}=\mbox{vec}({\cal E}_{t})$. Write
@@
$=A\Theta A^{\top}+\sum_{i,j=1}^{r}\frac{1}{T}\sum_{t=1}^{T}\left(f_{i,t}f_{j,t}-\mathbb{E}f_{i,t}f_{j,t}\right)a_{i}a_{j}^{\top}+\frac{1}{T}\sum_{t=1}^{T}e_{t}e_{t}^{\top}$
$\quad+\frac{1}{T}\sum_{t=1}^{T}\sum_{i=1}^{r}f_{i,t}a_{i}e_{t}^{\top}+\frac{1}{T}\sum_{t=1}^{T}\sum_{i=1}^{r}f_{i,t}e_{t}a_{i}^{\top}$
@@
As $1/\gamma=2/\gamma_{1}+1/\gamma_{2}$, by Theorem 1 in Merlevède et al., (2011),
$\mathbb{P}\left(T\left|u_{j}^{\top}\Theta^{-1/2}(\widehat{\Theta}-\Theta)\Theta^{-1/2}u_{j}\right|\geq x\right)$ $\leq T\exp\left(-\frac{x^{\gamma}}{c_{1}}\right)+\exp\left(-\frac{x^{2}}{c_{2}T}\right)$
@@
$\mathbb{P}\left(T\left\|\Theta^{-1/2}(\widehat{\Theta}-\Theta)\Theta^{-1/2}\right\|_{\rm 2}/3\geq x\right)$ $\leq N_{r,1/3}^{2}T\exp\left(-\frac{x^{\gamma}}{c_{1}}\right)+N_{r,1/3}^{2}\exp\left(-\frac{x^{2}}{c_{2}T}\right)$
@@
Choosing $x\asymp\sqrt{T(r+\log T)}+(r+\log T)^{1/\gamma}$, in an event $\Omega_{1}$ with probability at least $21-T^{-c_{1}}/$ 2,
$\left\|\Theta^{-1/2}(\widehat{\Theta}-\Theta)\Theta^{-1/2}\right\|_{\rm 2}\leq C\sqrt{\frac{r+\log(T)}{T}}+\frac{C(r+\log T)^{1/\gamma}}{T}.$
@@
$\|\Delta_{1}\|_{\rm 2}$ $\leq\|A\|_{\rm 2}^{2}\lambda_{1}\cdot\left\|\Theta^{-1/2}(\widehat{\Theta}-\Theta)\Theta^{-1/2}\right\|_{\rm 2}$
@@
Note that $\mathbb{E}\|e_{t}\|_{2}^{2}=\mathbb{E}\xi_{t}^{\top}H^{\top}H\xi_{t}=\mathbb{E}\mbox{tr}(H^{\top}H\xi_{t}\xi_{t}^{\top})=\mbox{tr}(H^{\top}H)=\mbox{tr}(HH^{\top})\asymp d$, and $\|H^{\top}H\|_{\rm F}^{2}=\|HH^{\top}\|_{\rm F}^{2}=\|\Sigma_{e}\|_{\rm F}^{2}\asymp d$. Choosing $x\asymp d$, we have
$\mathbb{P}\left(\|e_{t}\|_{2}\geq C\sqrt{d}\right)\leq 4\exp\left(-C^{\prime}d^{\frac{\vartheta}{2\vartheta+4}}\right).$
Let $N:=\|e_{t}e_{t}^{\top}\mathbf{1}_{\{\|e_{t}\|_{2}\leq C\sqrt{d}\}}\|_{2}$ and $\sigma_{0}^{2}:=\|\sum_{t=1}^{T}\mathbb{E}(e_{t}e_{t}^{\top}\mathbf{1}_{\{\|e_{t}\|_{2}\leq C\sqrt{d}\}})^{2}\|_{2}$. Then, by Assumption 4.1, $N\leq C^{2}d$ and
@@
$\mathbb{P}\left(\left\|\sum_{t=1}^{T}\left[e_{t}e_{t}^{\top}\mathbf{1}_{\{\|e_{t}\|_{2}\leq C\sqrt{d}\}}-\mathbb{E}e_{t}e_{t}^{\top}\mathbf{1}_{\{\|e_{t}\|_{2}\leq C\sqrt{d}\}}\right]\right\|_{2}\geq x\right)\leq 2d\exp\left(-\frac{x^{2}/2}{\sigma_{0}^{2}+Nx/3}\right).$ (41)
Choosing $x\asymp\sqrt{Td\log(d)}+d\log(d)$, with probability at least $1-d^{-c_{1}}$,
$\left\|\frac{1}{T}\sum_{t=1}^{T}e_{t}e_{t}^{\top}\mathbf{1}_{\{\|e_{t}\|_{2}\leq C\sqrt{d}\}}-\mathbb{E}e_{t}e_{t}^{\top}\mathbf{1}_{\{\|e_{t}\|_{2}\leq C\sqrt{d}\}}\right\|_{2}\leq C_{1}\sqrt{\frac{d\log(d)}{T}}+C_{1}\cdot\frac{d\log(d)}{T}$ (42)
Define $M:=\{1\leq t\leq T:\|e_{t}\|_{2}\geq C\sqrt{d}\}$. Since $\mathbf{1}_{\{\|e_{t}\|_{2}\geq C\sqrt{d}\}}$ are independent Bernoulli random variable and $\log(T)\leq d^{\vartheta/(2\vartheta+4)}$, we have
$\mathbb{E}|M|=T\mathbb{P}\left(\|e_{t}\|_{2}\geq C\sqrt{d}\right)\leq 4T\exp\left(-C^{\prime}d^{\frac{\vartheta}{2\vartheta+4}}\right)\leq T^{-c_{2}}.$
@@
$\mathbb{P}\left(\left\|\sum_{t=1}^{T}e_{t}e_{t}^{\top}\mathbf{1}_{\{\|e_{t}\|_{2}\geq C\sqrt{d}\}}\right\|_{2}\geq x\right)$ $\leq\mathbb{P}\left(|M|\max_{t}\|e_{t}\|_{2}^{2}\geq x\right)$
@@
$\leq\exp\left(-T^{c_{2}}\right)+\mathbb{P}\left(\max_{t}\|e_{t}\|_{2}^{2}\geq x/C\right)$
Choosing $x\asymp d$, we have, with probability at least $1-\exp(-T^{c_{2}})-T^{-c_{2}}$,
$\left\|\frac{1}{T}\sum_{t=1}^{T}e_{t}e_{t}^{\top}\mathbf{1}_{\{\|e_{t}\|_{2}\geq C\sqrt{d}\}}\right\|_{2}\leq C_{2}\cdot\frac{d}{T}.$ (44)
@@
$\mathbb{P}\left(\left\|\mathbb{E}e_{t}e_{t}^{\top}\mathbf{1}_{\{\|e_{t}\|_{2}\geq C\sqrt{d}\}}\right\|_{2}>0\right)=\mathbb{P}\left(\|e_{t}\|_{2}\geq C\sqrt{d}\right)\leq 4\exp\left(-C^{\prime}d^{\frac{\vartheta}{2\vartheta+4}}\right)\leq T^{-c_{3}}.$ (45)
@@
$\leq\left\|\frac{1}{T}\sum_{t=1}^{T}e_{t}e_{t}^{\top}\mathbf{1}_{\{\|e_{t}\|_{2}\leq C\sqrt{d}\}}-\mathbb{E}e_{t}e_{t}^{\top}\mathbf{1}_{\{\|e_{t}\|_{2}\leq C\sqrt{d}\}}\right\|_{2}+\left\|\frac{1}{T}\sum_{t=1}^{T}e_{t}e_{t}^{\top}\mathbf{1}_{\{\|e_{t}\|_{2}\geq C\sqrt{d}\}}\right\|_{2}$
$\quad+\left\|\mathbb{E}e_{t}e_{t}^{\top}\mathbf{1}_{\{\|e_{t}\|_{2}\geq C\sqrt{d}\}}\right\|_{2}+\|\Sigma_{e}\|_{2}$
$\leq C_{2}\sqrt{\frac{d\log(d)}{T}}+C_{2}\cdot\frac{d\log(d)}{T}+C_{2}$
Next, consider $\|\Delta_{3}\|_{2}$. Note that $\|\Delta_{4}\|_{2}$ is the same as $\|\Delta_{3}\|_{2}$. Let ¯$\overline{\mathbb{P}}(\cdot)=\mathbb{P}(\cdot|F_{1},...,F_{T})$ and $\overline{\mathbb{E}}(\cdot)=\mathbb{E}(\cdot|F_{1},...,F_{T})$ be the conditional probability and conditional expectation given the factor process, respectively. Similar to the derivation for $\|\Delta_{2}\|_{2}$, let
$N_{1}:=\left\|\sum_{i=1}^{r}f_{it}a_{i}e_{t}^{\top}\mathbf{1}_{\{\|e_{t}\|_{2}\leq C\sqrt{d}\}}\right\|_{2},$
$\sigma_{1}^{2}:=\max\left\{\left\|\sum_{t=1}^{T}\overline{\mathbb{E}}\sum_{i,j=1}^{r}f_{it}f_{jt}a_{i}e_{t}^{\top}e_{t}a_{j}^{\top}\mathbf{1}_{\{\|e_{t}\|_{2}\leq C\sqrt{d}\}}\right\|_{2},\left\|\sum_{t=1}^{T}\overline{\mathbb{E}}\sum_{i,j=1}^{r}f_{it}f_{jt}e_{t}a_{i}^{\top}a_{j}e_{t}^{\top}\mathbf{1}_{\{\|e_{t}\|_{2}\leq C\sqrt{d}\}}\right\|_{2}\right\}.$
@@
$N_{1}\leq C\sqrt{d}\left\|\sum_{i=1}^{r}f_{it}a_{i}\right\|_{2},$
$\sigma_{1}^{2}\leq C_{3}Td\max\left\{\left\|\frac{1}{T}\sum_{t=1}^{T}\sum_{i,j=1}^{r}f_{it}f_{jt}a_{i}a_{j}^{\top}\right\|_{2},\left\|\frac{1}{Td}\sum_{t=1}^{T}\sum_{i,j=1}^{r}f_{it}f_{jt}a_{i}^{\top}a_{j}\right\|_{2}\right\}:=\sigma_{2}^{2}.$

### v2

@@ 7 Conclusion
Modeling high-dimensional tensor time series has attracted growing attention, driven by the increasing availability of multidimensional datasets that go beyond the classical panel data structure. This paper considers matrix and tensor factor models with a CP low-rank structure, extending traditional vector factor models to higher-order settings. We develop ISO procedures based on contemporary covariance, preserving the tensor data structure. Theoretical properties such as the rate of convergence and limiting distributions are investigated under settings where each tensor dimension is comparable to or exceeds the number of observations, and the tensor rank may be either fixed or divergent.
Unlike auto-covariance-based estimation methods, our method explores information from contemporary covariance and can consistently estimate both loadings and factors even when observations are uncorrelated, settings where auto-covariance methods may fail. Additionally, we propose two generalized eigenvalue-ratio estimators for rank selection and justify their consistency. Extensive simulation studies highlight the merits of our method over existing methods. Furthermore, the empirical application to sorted portfolios demonstrates the practical relevance of our approach.
@@ Appendix A Proofs of Main Theorem
Part (i). Recall $\widehat{\Sigma}=T^{-1}\sum_{t=1}^{T}{\cal Y}_{t}\otimes{\cal Y}_{t}$, $\Theta=W(\mathbb{E}F_{t}F_{t}^{\top})W$, $W=\operatorname{diag}(w_{1},...,w_{r})$, $a_{i}=\mbox{vec}(a_{i1}\otimes a_{i2}\otimes\cdots\otimes a_{iK})$, $d=d_{1}d_{2}...d_{K}$. Let $e_{t}=\mbox{vec}({\cal E}_{t})$. Write
@@
Part (ii). For simplicity, consider the most extreme case where $\min\{\lambda_{i}-\lambda_{i+1},\lambda_{i}-\lambda_{i-1}\}\leq c\lambda_{r}$ for all $i$, with $\lambda_{0}=\infty,\lambda_{r+1}=0$, and $c$ is sufficiently small constant. In such cases, we need to employ Procedure 2 to the entire sample covariance tensor $\widehat{\Sigma}$. Let the eigenvalue ratio $\kappa:=\lambda_{1}/\lambda_{r}=O(r)$. Without loss of generality, assume $\Theta_{11}\geq\Theta_{22}\geq\cdots\geq\Theta_{rr}$ and $h=1$ in Procedure 2. In general, the statement in the theorem holds for number of initialization $L\geq Cd^{2}\vee Cdr^{2\kappa^{2}}$, where $a\vee b=\max\{a,b\}$. We prove the statements through induction on factor index $i$ starting from $i=1$ proceeding to $i=r$. By the induction hypothesis, we already have estimators such that
@@
and $R^{(0)}=\phi^{(0)}$ is defined in (19) in Theorem 4.1. We shall show that the bound in (37) also holds for $i$ on the same event. Note that when $i=1$, the statement “the bound in (37) holds for $1\leq j\leq i-1$ for all $1\leq k\leq K$” is an empty set. Thus, we focus on proving only the inductive step.
@@
$\geq\left\|\Theta_{ii}\prod_{k=1}^{K}(a_{ik}^{\top}\widetilde{a}_{\ell k})^{2}\right\|_{2}-\left\|\sum_{(j_{1},j_{2})\neq(i,i)}^{r}\Theta_{j_{1},j_{2}}\prod_{k=1}^{K}(a_{j_{1}k}^{\top}\widetilde{a}_{\ell k})(a_{j_{2}k}^{\top}\widetilde{a}_{\ell k})\right\|_{2}-\left\|\Psi\times_{k=1}^{2K}\widetilde{a}_{\ell k}\right\|_{2},$
where $\Psi$ is defined by unfolding $\Psi^{*}$ into a $d_{1}\times d_{2}\times\cdots\times d_{K}\times d_{1}\times d_{2}\times\cdots\times d_{K}$ tensor. Let $\psi_{i}=CR^{(0)}/\Theta_{ii}+\delta_{1}^{2}\kappa$. By (A) and the last part of the proof of Lemma A.2, as $\|\Psi^{*}\|_{2}/\lambda_{r}\leq\phi_{0}^{2}$ and $(1+\delta_{1})\prod_{k=2}^{K}(\delta_{k}+\psi_{i})w\leq\phi_{0}^{2}$, it follows that
@@
$\|\Psi^{*}\|_{\rm 2}$ $\leq C\lambda_{1}\left(\sqrt{\frac{r+\log T}{T}}+\frac{(r+\log T)^{1/\gamma}}{T}\right)+C\left(\sqrt{\frac{d\log(d)}{T}}+\frac{d\log(d)(\log T)^{\frac{2\vartheta+4}{\vartheta}}}{T}+1\right)$
$\qquad\quad+C\lambda_{1}^{1/2}\left(\sqrt{\frac{d\log(d)}{T}}+\frac{\sqrt{dr}\log(d)(\log T)^{1+\frac{2}{\vartheta}+\frac{1}{\gamma_{1}}}}{T}\right).$ (39)
@@
Let $\Upsilon_{0}=T^{-1}\sum_{t=1}^{T}\sum_{i,j=1}^{r}w_{i}w_{j}f_{it}f_{jt}a_{i}a_{j}^{\top}$, $\overline{\mathbb{E}}(\cdot)=\mathbb{E}(\cdot|f_{it},1\leq i\leq r,1\leq t\leq T)$. Define $e_{t}=\mbox{vec}({\cal E}_{t})$. Write
@@
$=A\Theta A^{\top}+\sum_{i,j=1}^{r}\frac{1}{T}\sum_{t=1}^{T}w_{i}w_{j}\left(f_{i,t}f_{j,t}-\mathbb{E}f_{i,t}f_{j,t}\right)a_{i}a_{j}^{\top}+\frac{1}{T}\sum_{t=1}^{T}e_{t}e_{t}^{\top}$
$\quad+\frac{1}{T}\sum_{t=1}^{T}\sum_{i=1}^{r}w_{i}f_{i,t}a_{i}e_{t}^{\top}+\frac{1}{T}\sum_{t=1}^{T}\sum_{i=1}^{r}w_{i}f_{i,t}e_{t}a_{i}^{\top}$
@@
As $1/\gamma=2/\gamma_{1}+1/\gamma_{2}$ and $W=\operatorname{diag}(w_{1},...,w_{r})$, by Theorem 1 in Merlevède et al., (2011),
$\mathbb{P}\left(T\left|u_{j}^{\top}W^{-1}(\widehat{\Theta}-\Theta)W^{-1}u_{j}\right|\geq x\right)$ $\leq T\exp\left(-\frac{x^{\gamma}}{c_{1}}\right)+\exp\left(-\frac{x^{2}}{c_{2}T}\right)$
@@
$\mathbb{P}\left(T\left\|W^{-1}(\widehat{\Theta}-\Theta)W^{-1}\right\|_{\rm 2}/3\geq x\right)$ $\leq N_{r,1/3}^{2}T\exp\left(-\frac{x^{\gamma}}{c_{1}}\right)+N_{r,1/3}^{2}\exp\left(-\frac{x^{2}}{c_{2}T}\right)$
@@
Choosing $x\asymp\sqrt{T(r+\log T)}+(r+\log T)^{1/\gamma}$, in an event $\Omega_{1}$ with probability at least $1-T^{-c_{1}}/2$,
$\left\|W^{-1}(\widehat{\Theta}-\Theta)W^{-1}\right\|_{\rm 2}\leq C\sqrt{\frac{r+\log(T)}{T}}+\frac{C(r+\log T)^{1/\gamma}}{T}.$
@@
$\|\Delta_{1}\|_{\rm 2}$ $\leq\|A\|_{\rm 2}^{2}\lambda_{1}\cdot\left\|W^{-1}(\widehat{\Theta}-\Theta)W^{-1}\right\|_{\rm 2}$
@@
Note that $\mathbb{E}\|e_{t}\|_{2}^{2}=\mathbb{E}\xi_{t}^{\top}H^{\top}H\xi_{t}=\mathbb{E}\mbox{tr}(H^{\top}H\xi_{t}\xi_{t}^{\top})=\mbox{tr}(H^{\top}H)=\mbox{tr}(HH^{\top})\asymp d$, and $\|H^{\top}H\|_{\rm F}^{2}=\|HH^{\top}\|_{\rm F}^{2}=\|\Sigma_{e}\|_{\rm F}^{2}\asymp d$. Choosing $x\asymp d(\log T)^{(2\vartheta+4)/\vartheta}$, we have
$\mathbb{P}\left(\|e_{t}\|_{2}\geq C\sqrt{d}(\log T)^{(\vartheta+2)/\vartheta}\right)\leq 4\exp\left(-C^{\prime}d^{\frac{\vartheta}{2\vartheta+4}}\log T\right).$
Let $N:=\|e_{t}e_{t}^{\top}\mathbf{1}_{\{\|e_{t}\|_{2}\leq C\sqrt{d}(\log T)^{(\vartheta+2)/\vartheta}\}}\|_{2}$ and $\sigma_{0}^{2}:=\|\sum_{t=1}^{T}\mathbb{E}(e_{t}e_{t}^{\top}\mathbf{1}_{\{\|e_{t}\|_{2}\leq C\sqrt{d}(\log T)^{(\vartheta+2)/\vartheta}\}})^{2}\|_{2}$. Then, by Assumption 4.1, $N\leq C^{2}d(\log T)^{(2\vartheta+4)/\vartheta}$ and
@@
$\mathbb{P}\left(\left\|\sum_{t=1}^{T}\left[e_{t}e_{t}^{\top}\mathbf{1}_{\{\|e_{t}\|_{2}\leq C\sqrt{d}(\log T)^{(\vartheta+2)/\vartheta}\}}-\mathbb{E}e_{t}e_{t}^{\top}\mathbf{1}_{\{\|e_{t}\|_{2}\leq C\sqrt{d}(\log T)^{(\vartheta+2)/\vartheta}\}}\right]\right\|_{2}\geq x\right)\leq 2d\exp\left(-\frac{x^{2}/2}{\sigma_{0}^{2}+Nx/3}\right).$ (42)
Choosing $x\asymp\sqrt{Td\log(d)}+d(\log d)(\log T)^{(2\vartheta+4)/\vartheta}$, with probability at least $1-d^{-c_{1}}$,
$\left\|\frac{1}{T}\sum_{t=1}^{T}e_{t}e_{t}^{\top}\mathbf{1}_{\{\|e_{t}\|_{2}\leq C\sqrt{d}(\log T)^{(\vartheta+2)/\vartheta}\}}-\mathbb{E}e_{t}e_{t}^{\top}\mathbf{1}_{\{\|e_{t}\|_{2}\leq C\sqrt{d}(\log T)^{(\vartheta+2)/\vartheta}\}}\right\|_{2}\leq C_{1}\sqrt{\frac{d\log(d)}{T}}+C_{1}\cdot\frac{d\log(d)(\log T)^{\frac{2\vartheta+4}{\vartheta}}}{T}.$ (43)
Define $M:=\{1\leq t\leq T:\|e_{t}\|_{2}\geq C\sqrt{d}(\log T)^{(\vartheta+2)/\vartheta}\}$. Since $\mathbf{1}_{\{\|e_{t}\|_{2}\geq C\sqrt{d}(\log T)^{(\vartheta+2)/\vartheta}\}}$ are independent Bernoulli random variable, we have
$\mathbb{E}|M|=T\mathbb{P}\left(\|e_{t}\|_{2}\geq C\sqrt{d}(\log T)^{(\vartheta+2)/\vartheta}\right)\leq 4T\exp\left(-C^{\prime}d^{\frac{\vartheta}{2\vartheta+4}}\log T\right)\leq T^{-c_{2}}.$
@@
$\mathbb{P}\left(\left\|\sum_{t=1}^{T}e_{t}e_{t}^{\top}\mathbf{1}_{\{\|e_{t}\|_{2}\geq C\sqrt{d}(\log T)^{(\vartheta+2)/\vartheta}\}}\right\|_{2}\geq x\right)$ $\leq\mathbb{P}\left(|M|\max_{t}\|e_{t}\|_{2}^{2}\geq x\right)$
@@
$\leq\exp\left(-T^{c_{2}}\right)+\mathbb{P}\left(\max_{t}\|e_{t}\|_{2}^{2}\geq x/C\right).$
Choosing $x\asymp d(\log T)^{(2\vartheta+4)/\vartheta}$, we have, with probability at least $1-\exp(-T^{c_{2}})-T^{-c_{2}}$,
$\left\|\frac{1}{T}\sum_{t=1}^{T}e_{t}e_{t}^{\top}\mathbf{1}_{\{\|e_{t}\|_{2}\geq C\sqrt{d}(\log T)^{(\vartheta+2)/\vartheta}\}}\right\|_{2}\leq C_{2}\cdot\frac{d(\log T)^{\frac{2\vartheta+4}{\vartheta}}}{T}.$ (45)
@@
$\mathbb{P}\left(\left\|\mathbb{E}e_{t}e_{t}^{\top}\mathbf{1}_{\{\|e_{t}\|_{2}\geq C\sqrt{d}(\log T)^{(\vartheta+2)/\vartheta}\}}\right\|_{2}>0\right)=\mathbb{P}\left(\|e_{t}\|_{2}\geq C\sqrt{d}(\log T)^{(\vartheta+2)/\vartheta}\right)\leq 4\exp\left(-C^{\prime}d^{\frac{\vartheta}{2\vartheta+4}}\log T\right)\leq T^{-c_{3}}.$ (46)
@@
$\leq\left\|\frac{1}{T}\sum_{t=1}^{T}e_{t}e_{t}^{\top}\mathbf{1}_{\{\|e_{t}\|_{2}\leq C\sqrt{d}(\log T)^{(\vartheta+2)/\vartheta}\}}-\mathbb{E}e_{t}e_{t}^{\top}\mathbf{1}_{\{\|e_{t}\|_{2}\leq C\sqrt{d}(\log T)^{(\vartheta+2)/\vartheta}\}}\right\|_{2}$
$\quad+\left\|\frac{1}{T}\sum_{t=1}^{T}e_{t}e_{t}^{\top}\mathbf{1}_{\{\|e_{t}\|_{2}\geq C\sqrt{d}(\log T)^{(\vartheta+2)/\vartheta}\}}\right\|_{2}+\left\|\mathbb{E}e_{t}e_{t}^{\top}\mathbf{1}_{\{\|e_{t}\|_{2}\geq C\sqrt{d}(\log T)^{(\vartheta+2)/\vartheta}\}}\right\|_{2}+\|\Sigma_{e}\|_{2}$
$\leq C_{2}\sqrt{\frac{d\log(d)}{T}}+C_{2}\cdot\frac{d\log(d)(\log T)^{\frac{2\vartheta+4}{\vartheta}}}{T}+C_{2}.$
Next, consider $\|\Delta_{3}\|_{2}$. Note that $\|\Delta_{4}\|_{2}$ is the same as $\|\Delta_{3}\|_{2}$. Let $\overline{\mathbb{P}}(\cdot)=\mathbb{P}(\cdot|F_{1},...,F_{T})$ and $\overline{\mathbb{E}}(\cdot)=\mathbb{E}(\cdot|F_{1},...,F_{T})$ be the conditional probability and conditional expectation given the factor process, respectively. Similar to the derivation for $\|\Delta_{2}\|_{2}$, let
$N_{1}:=\left\|\sum_{i=1}^{r}w_{i}f_{it}a_{i}e_{t}^{\top}\mathbf{1}_{\{\|e_{t}\|_{2}\leq C\sqrt{d}(\log T)^{1+\frac{2}{\vartheta}}\}}\right\|_{2},$
$\sigma_{1}^{2}:=\max\left\{\left\|\sum_{t=1}^{T}\overline{\mathbb{E}}\sum_{i,j=1}^{r}w_{i}w_{j}f_{it}f_{jt}a_{i}e_{t}^{\top}e_{t}a_{j}^{\top}\mathbf{1}_{\{\|e_{t}\|_{2}\leq C\sqrt{d}(\log T)^{1+\frac{2}{\vartheta}}\}}\right\|_{2},\right$.
$\qquad\qquad\left.\left\|\sum_{t=1}^{T}\overline{\mathbb{E}}\sum_{i,j=1}^{r}w_{i}w_{j}f_{it}f_{jt}e_{t}a_{i}^{\top}a_{j}e_{t}^{\top}\mathbf{1}_{\{\|e_{t}\|_{2}\leq C\sqrt{d}(\log T)^{1+\frac{2}{\vartheta}}\}}\right\|_{2}\right\}.$
@@
$N_{1}\leq C\sqrt{d}(\log T)^{1+\frac{2}{\vartheta}}\left\|\sum_{i=1}^{r}w_{i}f_{it}a_{i}\right\|_{2},$
$\sigma_{1}^{2}\leq C_{3}Td\max\left\{\left\|\frac{1}{T}\sum_{t=1}^{T}\sum_{i,j=1}^{r}w_{i}w_{j}f_{it}f_{jt}a_{i}a_{j}^{\top}\right\|_{2},\left\|\frac{1}{Td}\sum_{t=1}^{T}\sum_{i,j=1}^{r}w_{i}w_{j}f_{it}f_{jt}a_{i}^{\top}a_{j}\right\|_{2}\right\}:=\sigma_{2}^{2}.$